OpenStudy (anonymous):

Prove, in complete sentences, whether the left expression is equal to the right expression and discuss which property applies. 6(x − 2) 6x − 12

OpenStudy (anonymous):

of course it is>>

OpenStudy (anonymous):

6(x-2) = 6x - 12 6*x - 6*2 = 6x - 12 6x - 12 = 6x - 12

OpenStudy (anonymous):

Why would they ask for a sentence the math is enough to show that they are equal